# Online Vectorized HD Map Construction using Geometry

### v1

@@ 3.4 Euclidean Loss and Objectives
GeMap [Ours] R50 65.1 normal-$\uparrow$ 5.3 59.8 normal-$\uparrow$ 3.6 63.2 normal-$\uparrow$ 3.1 62.7 normal-$\uparrow$ 4.0 15.6
@@
GeMap [Ours] R50 69.8 normal-$\uparrow$ 1.5 67.1 71.4 normal-$\uparrow$ 1.7 69.4 normal-$\uparrow$ 0.7 13.3
HDMapNet [ICRA’22] [14] EB0 & PP 29.6 16.3 46.7 31.0 -
@@
MapTR [ICLR’23] [18] R50 & Sec 62.3 55.9 69.3 62.5 -
GeMap [Ours] R50 & Sec 66.3 normal-$\uparrow$ 4.0 62.2 normal-$\uparrow$ 6.3 71.1 normal-$\uparrow$ 1.8 66.5 normal-$\uparrow$ 4.0 6.8
Table 2: Following BeMapNet [31], we also compare our methods on the NuScenes dataset under different weather conditions.
[…]
@@
GeMap [Ours] R50 66.0 normal-$\uparrow$ 3.9 64.3 normal-$\uparrow$ 3.8 54.4 normal-$\uparrow$ 1.6 61.5 normal-$\uparrow$ 3.1 15.6
@@
GeMap [Ours] R50 73.1 normal-$\uparrow$ 5.8 71.0 normal-$\uparrow$ 3.5 59.3 normal-$\uparrow$ 2.7 67.8 normal-$\uparrow$ 4.0 13.3
3.5 Geometry-Decoupled Decoder
In the Transformer decoder, we obtain the $N\times N_{v}$ fused queries by adding $N$ instance queries to $N_{v}$ point queries, which is suggested in [18]. We denote input tokens by ${\bm{E}}\in\mathbb{R}^{N_{A}\times D}$, where $D$ is the feature dimension and $N_{A}=N\cdot N_{v}$ is the number of fused queries. Intuitively, each fused query corresponds to a point on a predicted map instance. Self-Attention (SA) is formulated by
$\small{\rm SA}({\bm{E}},{\bm{M}})={\rm Softmax}\left(\frac{({\bm{E}}{\bm{W}}^{q})({\bm{E}}{\bm{W}}^{k})^{\top}}{\sqrt{D_{k}}}\odot{\bm{M}}\right){\bm{E}}{\bm{W}}^{v}\,,$ (9)
Therefore, we present Geometry-Decoupled Attention (GDA). First, we multiply a binary mask ${\bm{M}}$ to the computed attention map so that the tokens of a map instance are aggregated according to the tokens within the same map instances only, which allows the model to adjust points’ positions according to learned shape geometry. Denote the index of map instance that the $i$-th token belongs to as $\mathcal{I}_{i}$. For example, tokens 1,2$\dots$,$N_{v}$ belong to the first instance so that $\mathcal{I}_{1},\mathcal{I}_{2},\dots,\mathcal{I}_{N_{v}}=1$. Tokens $N_{v}$+1, $N_{v}$+2, $\dots$, $2\cdot N_{v}$ belong to the second instance so that $\mathcal{I}_{N_{v}+1},\mathcal{I}_{N_{v}+2},\dots,\mathcal{I}_{2N_{v}}=2$. With this notation, the binary mask ${\bm{M}}$ can be simply constructed by
The second attention is expected to model relations between tokens of different map instances. The attention mask is given by
Table 3: Comparison on Argoverse 2 dataset. Our method demonstrates significant performance improvements over previous methods.
[…]
@@ 3.5 Geometry-Decoupled Decoder
GeMap [Ours] R50 67.6 normal-$\uparrow$ 31.5 59.3 normal-$\uparrow$ 21.0 64.7 normal-$\uparrow$ 25.5 63.9 normal-$\uparrow$ 26.0 13.5
@@
GeMap [Ours] R50 75.7 normal-$\uparrow$ 3.6 69.2 normal-$\uparrow$ 6.3 70.5 normal-$\uparrow$ 3.4 71.8 normal-$\uparrow$ 4.4 12.1
4 Experiments
4.1 Experimental Setups
Table 4: Ablation study on NuScenes datasets. We only train the model for 24 epochs and use simplified objectives (Equation 8).
[…]
@@ 4.1 Experimental Setups
Full 53.6 $\uparrow$ 4.1 49.2 normal-$\uparrow$ 4.5 54.8 $\uparrow$ 1.1 52.6 normal-$\uparrow$ 3.3
Figure 6: Visualization results. Instances hard to construct are highlighted in orange boxes. Scenario (a) depicts a complex triangular road boundary with an extended tail. Scenarios (b) and (c) depict pedestrian crossings that can only be partially observed. In scenario (d), the BEV map is tilted and lane markings are obscured by vehicles. These challenging cases indicate the superiority and robustness of GeMap.
Datasets. To evaluate our method, we conduct experiments on the NuScenes dataset [1], a widely recognized large-scale autonomous driving dataset that includes 1,000 scenes captured by six RGB cameras with a 360-degree field of view, and provides precise annotations from LiDAR point clouds for HD map construction. For comparability with prior research [14, 18], we focus on three static categories of map instances: pedestrian crossings, lane dividers, and road boundaries. With dataset splits provided by BeMapNet [31], we also evaluate our methods under three weather conditions: sunny, cloudy, and rainy. Additionally, we used the Argoverse 2 dataset [39] as another benchmark, which consists of approximately 108,000 frames, each providing images from seven cameras.
Metrics. We evaluate the performance of our method using the widely adopted metric of Average Precision (AP) [14, 18]. Specifically, we categorize a prediction as a True Positive if the Chamfer Distance between the predicted instance and its ground truth counterpart is less than a predefined threshold. For our experiments, we set these thresholds at 0.5, 1.0, and 1.5 meters.
Implementation Details. Our model leverages 8 NVIDIA RTX 3090 GPUs for training. We adopt AdamW [25] as the optimizer and utilize Cosine Annealing with a linear warm-up phase [24] as the learning rate scheduler. For more details on hyperparameters, please refer to supplementary materials. Furthermore, our model operates with two objective configurations: a simplified one that adheres to the losses defined in Equation 8, and a comprehensive one that extends to include depth estimation, as well as PV and BEV segmentation losses, as delineated in Equation 7.
@@ 4.2 Main Results
NuScenes. As delineated in Table 1, our method delivers a state-of-the-art performance, achieving a mean Average Precision (mAP) of 69.4% using a camera-only setup and the ResNet50 [10] backbone. Our approach particularly improves the precision for identifying lane dividers and road boundaries, with enhancements of +1.5% points and +1.7%, respectively. Importantly, these gains in performance do not come at the cost of efficiency, as our method sustains inference speeds on par with the previously established state-of-the-art, measured in frames per second (FPS). Also, our method achieves significantly improved performance with extra LiDAR inputs, which indicates the generalizability of our method to multi-modality settings. Moreover, we also compare performance under different weather conditions, as presented in Table 2. Our method exceeds previous works under sunny, cloudy, and rainy scenarios, which shows the potential of geometry to raise model robustness to varied weather conditions.
Argoverse 2. Extending our evaluation to the Argoverse 2 dataset, as presented in Table 3, our method also presents a new SOTA performance of 71.8% mAP, outperforming the highly advanced MapTRv2 model by +4.4%. This achievement not only demonstrates the effectiveness of our approach on a different dataset but also emphasizes the adaptability and precision of our algorithm in the evolving landscape of autonomous driving technologies.
@@ 4.3 Ablation Study
Ablation studies on the NuScenes dataset are carried out to evaluate the individual contributions of our method’s components and other vision backbones.
Components Ablation. The ablation experiments, detailed in Table 4, involved training the model for a reduced duration of 24 epochs with a simplified objective described by Equation 8, to facilitate efficient analysis. These experiments confirm the significant role of GDA, which alone increases mAP by +1.9%. Moreover, applying the full suite of components results in a further mAP enhancement of +1.4%. A notable discovery is the detrimental effect on model performance when Euclidean loss is applied without GDA, which leads to a 0.5% reduction in mAP. This reinforces our position that conventional self-attention mechanisms are insufficient for encoding a variety of geometry.
Figure 7: Hyperparameter experiments. (a) Best performance is reached with balanced Euclidean shape and relation losses. (b) Model performances degrade with too large or small $\lambda$.
Backbone Ablation. Additionally, we investigate 3 vision backbones, and the results are presented in Table 5. Our method achieves significantly improved performance with more advanced vision backbones. Specifically, VoVNetV2-99 [13] and Swin-Transformer Tiny [22] outperform the ResNet50 baseline by $+2.8\%$ and $+2.6\%$ on ${\rm mAP}$.
Table 5: Ablation of vision backbones on Nuscenes datasets.
[…]
R50 69.8 67.1 71.4 69.4 13.3
Swin-T 72.8 70.4 72.8 72.0 10.0
V2-99 72.5 70.7 73.5 72.2 9.5
@@ 4.4 Visualization Analysis
More Accurate Shape Perception. Scenario (a) in Figure 6 demonstrates our method’s capability in accurately identifying complex road boundary shapes, such as an elongated triangular tail—typical in lane merging or diversion areas. This precise result is attributed to the shape geometry we propose, which allows for a more accurate description of the tail feature, in contrast to the MapTR model, which does not discern it as effectively.
Alleviating Occlusion Issues. The scenarios depicted in (b) and (c) of Figure 6 demonstrate the utility of our shape geometry in alleviating challenges associated with occlusions in partially visible instances. In particular, scenario (b) presents a pedestrian crossing whose segment is concealed, necessitating models to hypothesize the hidden geometry based on contextual semantic information. Our method leverages geometry between road boundaries and pedestrian crossings to reconstruct a more faithful representation of the occluded area. Scenario (c) features a black car that obscures part of a pedestrian crossing; nonetheless, our approach successfully deduces the overall structure of the crossing within a complex shape. Contrastingly, MapTR [18] struggles with such complex shape recovery under similar occlusion conditions.
Enhanced Robustness to Rotational Transformations. Scenario (d) in Figure 6 exemplifies the resilience of our method to rotational transformations, as evidenced when the ego-vehicle executes a right turn causing the BEV map to appear highly tilted. Additionally, this scenario features lanes that are extensively obscured by vehicular traffic. Despite these challenges, our approach more adeptly maintains the integrity of lane width and parallelism, which are key geometry, in contrast to the MapTR [18] model. This underscores the superior robustness of our geometric constructs against rotational distortions.
@@ 4.5 Hyperparameter Experiments
Comparative Impact of Euclidean Loss Components. Figure LABEL:fig:shape_rel shows that our method exhibits larger sensitivity to the Euclidean shape loss than to the relation loss, as evidenced by the steeper change of the performance curve. The optimal result is obtained when the shape and relation losses are relatively balanced by weights, implying that both contribute comparably to the model’s optimization process.
Optimization of Euclidean Loss Weighting. The performance of the model, as demonstrated in Figure LABEL:fig:lossw, peaks when the weight of the Euclidean loss, $\lambda$, is set to $5\times{10}^{-3}$ Notably, model performance degrades with a too large or too small $\lambda$, which informed our decision to use a $\lambda$ value of $5\times 10^{-3}$ for our experiments.
@@ 5 Conclusion
In this paper, we realize significant shape and relation geometry inherent in HD map instances and propose the GeMap. GeMap includes the integration of Euclidean shape and relation losses for auxiliary supervision. To further refine the model’s awareness of diverse geometry, we introduce a geometry-decoupled attention mechanism. GeMap has achieved state-of-the-art performances on both the NuScenes and Argoverse 2 datasets, underscoring its effectiveness. Despite these promising results, the current application of geometry remains fundamental, and future research could focus on more sophisticated representations or enhanced geometric patterns. Furthermore, the application of geometry extends beyond HD map construction, offering potential solutions to occlusion challenges in other autonomous driving tasks. We anticipate that these findings will inspire further research. Acknowledgments. This work was funded by National Natural Science Foundation of China (No. 62272045).
[…]
@@
  • More details on the method design (§ A).
  • Further quantitative experimental results (§ B).
  • Additional visualization results under three weather conditions (§ C).
Appendix A Additional Details
A.1 Objective Functions
Two Objective Configurations. As detailed in § 3.4, our method employs two distinct objective functions. The full objective function is defined as follows:

### v2

@@ 4.1 Experimental Setups
MapVR [NeurIPS’23] [47] R50 61.8 55.0 59.4 58.8 19.8
GeMap [Ours] R50 65.1 $\uparrow$ 3.3 59.8 $\uparrow$ 4.8 63.2 $\uparrow$ 3.8 62.7 $\uparrow$ 3.9 19.0
@@
GeMap [Ours] R50 69.8 $\uparrow$ 1.5 67.1 71.4 $\uparrow$ 1.7 69.4 $\uparrow$ 0.7 15.8
HDMapNet [ICRA’22] [16] EB0 & PP 29.6 16.3 46.7 31.0 0.7
@@
MapTR [ICLR’23] [20] R50 & Sec 62.3 55.9 69.3 62.5 6.7
MapVR [NeurIPS’23] [47] R50 & PP 62.7 60.4 67.2 63.5 -
GeMap [Ours] R50 & Sec 66.3 $\uparrow$ 3.6 62.2 $\uparrow$ 1.8 71.1 $\uparrow$ 3.9 66.5 $\uparrow$ 3.0 7.3
MapTRv2 [Arxiv’23] [21] R50 & Sec 65.6 66.5 74.8 69.0 6.6
GeMap [Ours] R50 & Sec 69.8 $\uparrow$ 4.2 68.0 $\uparrow$ 1.5 73.4 70.4 $\uparrow$ 1.4 6.8
MapTRv2 [Arxiv’23] [21] V2-99 73.7 71.4 75.0 73.4 10.1
GeMap [Ours] Swin-T 72.8 70.4 72.8 72.0 11.4
GeMap [Ours] V2-99 76.0 $\uparrow$ 2.3 74.3 $\uparrow$ 2.9 77.7 $\uparrow$ 2.7 76.0 $\uparrow$ 2.6 10.8
Table 2: Following BeMapNet [34], we also compare GeMap on the nuScenes dataset under different weather conditions.
Methods Backbone Dns. Loss mAPsun()($\uparrow$)( ↑ ) mAPcld()($\uparrow$)( ↑ ) mAPrny()($\uparrow$)( ↑ ) mAPavg()($\uparrow$)( ↑ ) FPS()($\uparrow$)( ↑ )
@@
GeMap [Ours] R50 66.0 $\uparrow$ 3.9 64.3 $\uparrow$ 3.8 54.4 $\uparrow$ 1.6 61.5 $\uparrow$ 3.1 19.0
@@
GeMap [Ours] R50 73.1 $\uparrow$ 5.8 71.0 $\uparrow$ 3.5 59.3 $\uparrow$ 2.7 67.8 $\uparrow$ 4.0 15.8
Table 3: Comparison on Argoverse 2 dataset. GeMap demonstrates significant performance improvements over previous methods.
Methods Backbone Dns. Loss APdiv()($\uparrow$)( ↑ ) APped()($\uparrow$)( ↑ ) APbnd()($\uparrow$)( ↑ ) mAP()($\uparrow$)( ↑ ) FPS()($\uparrow$)( ↑ )
@@
GeMap [Ours] R50 67.6 $\uparrow$ 31.5 59.3 $\uparrow$ 21.0 64.7 $\uparrow$ 25.5 63.9 $\uparrow$ 26.0 16.7
@@
GeMap [Ours] R50 75.7 $\uparrow$ 3.6 69.2 $\uparrow$ 6.3 70.5 $\uparrow$ 3.4 71.8 $\uparrow$ 4.4 13.8
Table 4: Ablation study on the nuScenes dataset. We train the model for 24 epochs. * denotes replacing GDA with 2 layers of vanilla self-attention.
Method $\mathcal{L}_{\rm shp}$ (§ 3.3.1) $\mathcal{L}_{\rm rel}$ (§ 3.3.2) GDA (§ 3.5) $\mathrm{AP}_{div}(\uparrow)$ $\mathrm{AP}_{ped}(\uparrow)$ $\mathrm{AP}_{bnd}(\uparrow)$ $\mathrm{mAP}(\uparrow)$
@@
Replace GDA with 2-SA 2-SA 53.5 46.2 54.4 51.4
Full 53.6 $\uparrow$ 4.1 49.2 $\uparrow$ 4.5 54.8 $\uparrow$ 1.1 52.6 $\uparrow$ 3.3
Datasets. To evaluate GeMap, we conduct experiments on the nuScenes dataset [1], a widely adopted large-scale autonomous driving dataset that includes 1,000 scenes captured by six RGB cameras with a 360-degree field of view, and provides precise annotations from LiDAR point clouds for HD map construction. For comparability with prior research [16, 20], we focus on three static categories of map instances: pedestrian crossings, lane dividers, and road boundaries. With dataset splits provided by BeMapNet [34], we also evaluate GeMaps under three weather conditions: sunny, cloudy, and rainy. Additionally, we use the Argoverse 2 dataset [42] as another benchmark, which consists of approximately 108,000 frames, each providing images from seven cameras.
Metrics. We evaluate the performance of GeMap using the widely adopted metric of Average Precision (AP) [16, 20]. Specifically, we categorize a prediction as a True Positive if the Chamfer Distance between the predicted instance and its ground truth counterpart is less than a predefined threshold. For our experiments, we set these thresholds at 0.5, 1.0, and 1.5 meters.
Implementation Details. GeMap leverages 8 NVIDIA RTX 3090 GPUs for training. We adopt AdamW [28] as the optimizer and utilize Cosine Annealing with a linear warm-up phase [27] as the learning rate scheduler. For more details on hyperparameters, please refer to appendix § 0.A.2.
@@ 4.2 Main Results
Results on nuScenes. As delineated in Table 1, GeMap delivers a state-of-the-art performance, achieving a mean Average Precision (mAP) of 69.4% using the camera-only setup and ResNet50 [10] backbone. GeMap particularly improves the precision for identifying dividers and boundaries, with enhancements of +1.5% and +1.7%, respectively. Importantly, these performance gains do not come at a high cost of efficiency, as GeMap sustains inference speeds on par with and even exceeds the previously established state-of-the-art, measured in frames per second (FPS). With more powerful vision backbones, GeMap can achieve significantly improved performance. Specifically, VoVNetV2-99 [14] and Swin-T [25] outperform the ResNet50 baseline by $+6.6\%$ and $+2.6\%$ in ${\rm mAP}$, respectively. Also, GeMap achieves significantly improved performance with extra LiDAR inputs, which indicates the generalizability of GeMap to multi-modality settings. Moreover, we also compare performance under different weather conditions in Table 2. GeMap exceeds previous works under sunny, cloudy, and rainy scenarios, which shows the potential of geometry to raise model robustness to varied weather conditions.
Results on Argoverse 2. Extending our evaluation to the Argoverse 2 dataset, as presented in Table 3, GeMap also presents a new SOTA performance of 71.8% mAP, outperforming the highly advanced MapTRv2 model by +4.4%. This achievement not only demonstrates the effectiveness of GeMap on a different dataset but also emphasizes the adaptability and precision of GeMap in the evolving landscape of autonomous driving technologies.
@@ 4.3 Ablation Study
Ablation studies on the nuScenes dataset are carried out to evaluate the individual contributions of GeMap’s components and other vision backbones.
Components Ablation. The ablation experiments, detailed in Table 4, involved training the model for a reduced duration of 24 epochs without dense prediction losses, to facilitate efficient analysis. These experiments confirm the significant role of GDA, which alone increases mAP by +1.9%. Moreover, applying the full suite of components results in a further mAP enhancement of +1.4%. A notable discovery is the detrimental effect on model performance when Euclidean Loss is applied without GDA, which leads to a 0.5% reduction in mAP. This reinforces our position that conventional self-attention mechanisms are insufficient for encoding a variety of geometric properties.
Figure 6: Visualization results. Instances that are hard to construct are highlighted in orange boxes. Scenario (a) depicts a complex triangular road boundary. Scenario (b) includes a divider that is hard to recognize according to strong sunlight. Scenario (c) depicts pedestrian crossings that can only be partially observed. In scenario (d), the BEV map is tilted and lane markings are obscured by vehicles. These challenging cases indicate the superiority and robustness of GeMap.
Effectiveness of GDA. In the proposed GDA (3.5), we decouple shape and relation learning, sequentially applying Euclidean shape and relation attention. In addition, we also attempt to intuitively double the self-attention layers in one block, referred to as “2-SA”. As shown in Table 4, GDA outperforms vanilla self-attention by +1.2% mAP, indicating the superiority of GDA when combined with Euclidean Loss.
Figure 7: Hyperparameter experiments. (a) The best performance is reached with balanced Euclidean shape loss and relation loss. (b) Model performances degrade with too large or too small $\lambda$.
@@ 4.4 Visualization Analysis
More Precise Shape Awareness. Scenario (a) in Figure 6 demonstrates GeMap’s capability in accurately identifying complex road boundary shapes, such as triangular boundaries. This precise result is attributed to the shape geometry we propose, which allows for an inherent description of the triangle feature, in contrast to baselines which do not discern it as effectively. A more irregularly shaped pedestrian crossing can be viewed in scenario (c) of Figure 6 and GeMap provides more precise construction than other methods.
Better Relation Awareness. With the help of Euclidean Relation Clues, GeMap can better infer parallelism and perpendicular in scenario (a) of Figure 6. More interestingly, in scenario (b), the strong sunlight makes the divider highlighted hard to recognize. However, according to the understanding of lane width patterns, GeMap has the sense that there should be a divider.
Alleviating Occlusion Issues. The scenarios depicted (c) of Figure 6 demonstrate the utility of our shape geometry in alleviating challenges associated with occlusions in partially visible instances. In particular, scenario (c) features a black car that obscures part of a pedestrian crossing; nonetheless, GeMap successfully deduces the overall structure of the crossing within a complex shape. In contrast, these baseline models [20, 21, 4] struggle with such complex shape recovery under similar occlusion conditions.
Enhanced Robustness to Rotational Transformations. Scenario (d) in Figure 6 exemplifies the resilience of GeMap to rotational transformations, as evidenced when the ego-vehicle executes a right turn causing the BEV map to appear highly tilted. Additionally, this scenario features lanes that are extensively obscured by vehicular traffic. Despite these challenges, GeMap more adeptly maintains the integrity of lane width and parallelism, which are key geometry, in contrast to the baseline model. This underscores the superior robustness of our geometric constructs against rotational distortions.
@@ 4.5 Hyperparameter Experiments
Comparative Impact of Euclidean Losses. Figure LABEL:fig:shape_rel shows that GeMap exhibits larger sensitivity to the shape loss than to the relation loss, as evidenced by the steeper change of the performance curve. The optimal result is obtained when the shape and relation losses are relatively balanced by weights, implying that both contribute comparably to the model’s optimization process.
Optimization of Euclidean Loss Weighting. The performance of the model, as demonstrated in Figure LABEL:fig:lossw, peaks when the weight of the Euclidean Loss, $\lambda$, is set to $5\times{10}^{-3}$ Notably, model performance degrades with a too large or too small $\lambda$, which informed us to use a $\lambda$ value of $5\times 10^{-3}$ for our experiments.
Distance Weighting. Experimental results of distance weighting are presented in Table 5. We change the order of distance ($p$) in Equation 5 and $p\in\{1,2,4\}$ corresponds to “Linear”, “Square” and “4th Power” respectively. Moreover, we also try to treat instance pairs equally and it delivers the best performance.
Table 5: Distance weighting on the nuScenes dataset.
Strategy $\text{AP}_{div}(\uparrow)$ $\text{AP}_{ped}(\uparrow)$ $\text{AP}_{bnd}(\uparrow)$ mAP()($\uparrow$)( ↑ )
Linear 53.5 46.6 56.2 52.1
Square 53.2 47.9 55.2 52.1
4th Power 53.2 47.4 55.8 52.1
Equal 53.6 49.2 54.8 52.6
@@ 5 Conclusion
In this paper, we realize significant shape and relation geometry inherent in HD map instances and propose the GeMap. GeMap includes the integration of Euclidean shape and relation losses for auxiliary supervision. To further refine the model’s awareness of diverse geometry, we introduce the Geometry-Decoupled Attention mechanism. GeMap has achieved state-of-the-art performances on both the nuScenes and Argoverse 2 datasets, underscoring its effectiveness. Despite these promising results, the current application of geometry remains fundamental, and future research could focus on more sophisticated representations or enhanced geometric patterns. Furthermore, the application of geometry extends beyond HD map construction, offering potential solutions to occlusion challenges in other autonomous driving tasks. We anticipate that these findings will inspire further research.
Acknowledgments. This work is partially supported by the National Natural Science Foundation of China (No. 62272045, No. 8326014), The Shun Hing Institute of Advanced Engineering (No. 8115074), and CUHK Direct Grants (No. 4055190).
@@
  • More details on the method design (§ 0.A).
  • Further quantitative experimental results (§ 0.B).
  • Additional visualization results under three weather conditions (§ 0.C).
Appendix 0.A Additional Details
0.A.1 Objective Functions
Objective Configurations. Our method employs two distinct objective functions. The full objective function is defined as follows: